If random variable (x) takes vales 1, 2, 3, 4 such

that 2 P(x=1) = 3 P (x=2) = P(x=3) = 5 P (x=4)

Find the probability distribution function (PDF) and cumulative distribution function

(CDF), also state the properties of PDF and CDF.


1
Expert's answer
2021-02-24T06:08:38-0500

2 P(x=1) = 3 P (x=2) = P(x=3) = 5 P (x=4) =t

Then P(x=1)=t/2, P (x=2)=t/3, P(x=3)=t, P (x=4) =t/5

1=P(x=1) + P(x=2) + P(x=3)+ P(x=4) = t/2+t/3+t+t/5 = 61t/30

Therefore, t=30/61

P(x=1) = 15/61

P(x=2) = 10/61

P(x=3) = 30/61

P(x=4) = 6/61

Probability distribution function (PDF) is

"P(x=t)=\begin{Bmatrix}
 15/61 & if\ t=1 \\
 10/61 & if\ t=2 \\
 30/61 & if\ t=3 \\
 6/61 & if\ t=4 \\
0 & else
\end{Bmatrix}"

Its properties:

1) This PDF takes non-zero values in a finite set of points (discreteness).

2) All values of PDF are non-negative real numbers, not greater than 1.

3) The sum of all non-zero values of PDF equals 1.


Cumulative distribution function (CDF) is "P(x\leq t)".

If t<1 then CDF(t)=0, because of x does not take any values t<1.

If "1\leq t<2" then "P(x\leq t)=P(x=1)=15/61"

If "2\leq t<3" then "P(x\leq t)=P(x=1 \vee x=2)=P(x=1)+P(x=2)=15/61+10/61=25/61"

If "3\leq t<4" then "P(x\leq t)=P(x=1)+P(x=2)+P(x=3)=15/61+10/61+30/61=55/61"

If "t>4" then "x\leq t" in any case and"P(x\leq t)=1".

Finally,

"P(x\leq t)=\begin{Bmatrix}
0 & if\ t<1 \\
 15/61 & if\ 1\leq t<2 \\
 25/61 & if\ 2\leq t<3 \\
 55/61 & if\ 3\leq t<4 \\
 1 & if\ t\geq 4
\end{Bmatrix}"

Its properties:

1) CDF takes its values in the segment [0,1].

2) CDF is monotoneously increases everywhere

3) CDF has limits: "\lim\limits_{t\to-\infty}P(x\leq t)=0" and "\lim\limits_{t\to+\infty}P(x\leq t)=1".

4) CDF is continuous at every point t, such that PDF(t)=0.